OpenStudy (anonymous):

Two lines, A and B, are represented by the equations given below: Line A: 2x + 2y = 8 Line B: x + y = 4 Which statement is true about the solution to the set of equations?

OpenStudy (anonymous):

A.it is (8.4) B.it is (4,8) C.There is no solution D. Infinitely many solutions

pooja195 (pooja195):

2x + 2y = 8<---e1 x + y = 4 <---e2 Start by isolating x by subtracting it y=4+x now that we know the value of y we simply plug it into e1 2x+2(4+x)=8 solve for x

OpenStudy (anonymous):

so it will be 0?

pooja195 (pooja195):

Correct :) Now take x=0 and plug it into e2 y=4+0 4+0=?

OpenStudy (anonymous):

4 so will it be D?

pooja195 (pooja195):

I messed up x+y=4 for x: x+y+−y=4+−y(Add -y to both sides) x=−y+4 Step: Substitute−y+4 fo rx in 2x+2y=8: 2x+2y=8 2(−y+4)+2y=8 8=8(Simplify both sides of the equation) 8+−8=8+−8(Add -8 to both sides) 0=0 Therefore D is indeed correct :-)

OpenStudy (anonymous):

Thank you so so so much!!!!!!!!!!!!!

pooja195 (pooja195):

yw ^_^